OpenStudy (thequestion):

Solve, using the substitution method. y – 2x = 8 16 + 4x = 2y A. The solution is (1, 10). B. The solution is (24, 0). C. There is no solution. D. There are an infinite number of solutions.

OpenStudy (anonymous):

y-2x=8 y=8+2x sustitute it in other equation

OpenStudy (anonymous):

16+4x=2*(8+2x)=16+4x as both side are equal ...watever value of x u will take it will alwyz satisfy